OpenStudy (anonymous):

HELP!!! find all the real zeros of the polynomial function and state their multiplicities P P(x)=x^4-2x^2+8x-8

OpenStudy (anonymous):

Try to factor first.

OpenStudy (anonymous):

I got -1\[\pm \sqrt{-7} /2\]

OpenStudy (anonymous):

for the 1st one

OpenStudy (gorv):

x^4-5x^2+4 =x^4-4x^2-x^2+4=0 x^2(x^2-4)-1(x^2-4)=0 (x^2-4)(x^2-1)=0

OpenStudy (anonymous):